# LEASE: Offline Preference-based Reinforcement Learning with High Sample Efficiency

### v1

@@ A-C Proof of Theorem 2
Before proof, we give the related Assumption and Lemmas for Theorem 2. Firstly, we assume the the reward class $\mathcal{R}$ is realizable (Assumption 1) and bounded (Assumption 3).
Assumption 3 (Boundedness).
@@
The proof of Proposition 3 can be found in Appendix A-D. Note that, the above reward model is only trained from offline preference dataset. Through generating more preference data, the upper bound can be tighter. However, the generated data may be inaccuracy, which may improve the reward gap. The selecting mechanism can effectively solve this problem.
The Proof for Theorem 2: This Theorem aims to give the lower bound of term $J(\widehat{\pi},R^{*})-J(\mu,R^{*})$. We prove this from ideal case and actual case.
1) Ideal case: There is no need to consider distribution shift and empirical error problem. The policy $\widehat{\pi}$ is directly learned by $\widehat{\pi}=\max_{\pi}J(\pi,\widehat{R})$. Thus, $J(\widehat{\pi},\widehat{R})\geq J(\mu,\widehat{R})$ holds. Then, we can derive
@@
$\leq$ $\big{|}J(\mu,R^{*})-J(\mu,\widehat{R})\big{|}+\big{|}J(\widehat{\pi},R^{*})-J(\widehat{\pi},\widehat{R})\big{|}.$
Since $J(\pi,R):=\mathbb{E}_{(s,a)\sim d_{T}^{\pi}(s,a)}[R(s,a)]/(1-\gamma)$, $|J(\pi,R^{*})-J(\pi,\widehat{R})|$ can be written as
@@
$\Big{|}J(\pi,R^{*})-J(\pi,\widehat{R})\Big{|}\leq\frac{\mathscr{C}_{\mathcal{R}}(\pi)}{1-\gamma}\Big{|}\mathbb{E}_{(s,a)\sim d_{T}^{\pi}}\big{[}R^{*}(s,a)-\widehat{R}(s,a)\big{]}\Big{|}.$
@@
2) Actual case: Offline RL suffers from distribution shift problem and it is a necessity to consider empirical error. Therefore, many offline RL algorithms incorporate conservatism into policy to overcome distribution shift, that is learning policy through $\widehat{\pi}=\max_{\pi}J(\pi,\widehat{R})-P(\pi)$. Therefore, $J(\widehat{\pi},\widehat{R})\geq J(\mu,\widehat{R})-\xi$ instead of $J(\widehat{\pi},\widehat{R})\geq J(\mu,\widehat{R})$ in actual implantation. The term $\xi$ depends on the algorithm itself. Here, we take CQL as a example. The policy gap is proven in Theorem 3.6 of [4]. The detailed is given as follow.
@@
where ||$\cdot$|| ⋅ | denotes cardinality of a specific set, $D(s)=\sum_{a}\widehat{\pi}(a|s)(\frac{\widehat{\pi}(a|s)}{\mu(a|s)}-1)$ and $C_{T,R}$ is the empirical coefficient. It consists of two terms: the first term $\xi_{1}$ captures the decrease in policy performance due to sampling error. The second term $\xi_{2}$ captures the increase in policy performance in empirical setting [4].
Therefore, when considering distribution shift and empirical error, we can derive
@@ A-D Proof of Proposition 3
where $y\in\{0,1\}$, $R^{*}$ is the true reward model and $\widehat{R}$ is the learned reward model. $P_{R}(0~{}|~{}\sigma_{0},\sigma_{1})$ is the probability that $\sigma_{0}$ is preferable $\sigma_{1}$ under reward model $R$. $\mathcal{N}_{\mathcal{R}}(1/N)$ is the bracketing number defined in Definition 3. This proof can be found in Lemma 2 of previous work [18] and Proposition 14 of [23].
@@
Then, combing Eqs. (A.12) and (A.14), we can further derive
@@
where $[a,b]$ is the range of values that each $X_{i}$ can take. Then, for Eq. (A.15), we set $X_{i}$ as $R^{*}(s_{j},a_{j})-R(s_{j},a_{j})$ and $X_{i}\in[-2R_{max},2R_{max}])$. Then, the below equation holds

### v2

@@ A-C Proof of Theorem 2
The proof of Proposition 3 can be found in Appendix A-D. Note that the above reward model is only trained from the offline preference dataset. Through generating more preference data, the upper bound can be tighter. However, the generated data may be inaccurate, which may improve the reward gap. The selection mechanism can effectively solve this problem.
The Proof for Theorem 2: This Theorem aims to give the lower bound of term $J(\widehat{\pi},R^{*})-J(\mu,R^{*})$. We prove this from the ideal case and the actual case.
[…]
@@
[…]
@@
$\Big|J(\pi,R^{*})-J(\pi,\widehat{R})\Big|=\frac{\mathscr{C}_{\mathcal{R}}(\pi)}{1-\gamma}\Big|\mathbb{E}_{(s,a)\sim d_{T}^{\mu 0}}\big[R^{*}(s,a)-\widehat{R}(s,a)\big]\Big|.$
@@
[…]
@@
where $|\cdot|$ denotes cardinality of a specific set, $D(s)=\sum_{a}\widehat{\pi}(a|s)(\frac{\widehat{\pi}(a|s)}{\mu(a|s)}-1)$ and $C_{T,R}$ is the empirical coefficient. It consists of two terms: the first term $\xi_{1}$ captures the decrease in policy performance due to sampling error. The second term $\xi_{2}$ captures the increase in policy performance in the empirical setting [4]. Therefore, when considering distribution shift and empirical error, we can derive
@@ A-D Proof of Proposition 3
where $y\in\{0,1\}$, $R^{*}$ is the true reward model and $\widehat{R}$ is the learned reward model. $P_{R}(0~|~\sigma_{0},\sigma_{1})$ is the probability that $\sigma_{0}$ is preferable $\sigma_{1}$ under reward model $R$. $\mathcal{N}_{\mathcal{R}}(1/N)$ is the bracketing number defined in Definition 3. This proof can be found in Lemma 2 of the previous work [18] and Proposition 14 of [23].
@@
Then, combining Eqs. (A.12) and (A.14), we can further derive
@@
where $[a,b]$ is the range of values that each $X_{i}$ can take. Then, for Eq. (A.15), we set $X_{i}$ as $R^{*}(s_{j},a_{j})-R(s_{j},a_{j})$ and $X_{i}\in[-2R_{max},2R_{max}])$. Then, the following equation holds.
@@
Figure A.1: The comparison between prediction value by the learned rewards and their ground truths for different methods under (a) pen-expert and (b) hammer-expert datasets.
TABLE A.1: Comparison results between URLHF using fewer preference data and LEASE, where the offline RL algorithm is CQL and URLHF using fewer data is denoted as $\texttt{URLHF}^{*}$.
Task Name $\textbf{URLHF}^{*}$ URLHF FEWER LEASE
walker2d-m $76.1\pm 0.8$ $6.0\pm 0.9 77$$7.4\pm 0.6$ $\mathbf{78.4\pm 0.9}$
walker2d-m-e $86.8\pm 18.6$ $92.8\pm 22.4$ $77.7\pm 0.3$ $\mathbf{98.6\pm 18.1}$
hopper-m $\mathbf{56.6\pm 2.4}$ $4.7\pm 3.4 55$$5.8\pm 2.8$ $\mathbf{56.5\pm 0.6}$
hopper-m-e $55.3\pm 0.9$ $\mathbf{57.4\pm 4.9}$ $53.6\pm 0.9$ $56.4\pm 0.8$
halfcheetah-m $43.3\pm 0.2$ $\mathbf{43.4\pm 0.1}$ $\mathbf{43.5\pm 0.1}$ $\mathbf{43.5\pm 0.1}$
halfcheetah-m-e $58.9\pm 2.3$ $\mathbf{62.7\pm 7.1}$ $48.3\pm 0.7$ $53.2\pm 3.1$
Mujoco Average $62.8\pm 4.2$ $\mathbf{64.5\pm 6.5}$ $59.4\pm 0.9$ $\mathbf{64.4\pm 4.0}$
pen-human $\mathbf{17.7\pm 13.0}$ $9.8\pm 14.1$ $0.5\pm 3.0$ $3.8\pm 4.6$
pen-expert $114.6\pm 53.7$ $\mathbf{138.3\pm 5.2}$ $128.1\pm 0.7$ $130.6\pm 1.8$
door-human $1.7\pm 1.1$ $\mathbf{4.7\pm 5.9}$ $0.2\pm 1.0$ $\mathbf{4.7\pm 8.8}$
door-expert $\mathbf{103.3\pm 0.5}$ $\mathbf{103.9\pm 0.8}$ $103.0\pm 0.9$ $\mathbf{103.2\pm 0.7}$
hammer-human $0.7\pm 0.1$ $\mathbf{0.9\pm 0.3}$ $0.3\pm 0.0$ $0.3\pm 0.0$
hammer-expert $117.4\pm 2.7$ $120.2\pm 6.8$ $124.1\pm 2.1$ $\mathbf{126.3\pm 1.2}$
Adroit Average $59.2\pm 11.8$ $\mathbf{63.0\pm 5.5}$ $59.4\pm 1.3$ $61.5\pm 2.9$
A-E Other Experimental Results
These part presents supplementary experimental results, including: 1) the performance evaluation of the reward model on other tasks, 2) a comparative analysis between the proposed method LEASE and URLHF when trained on identical datasets, 3) an investigation of our framework’s effectiveness in model-based offline RL settings, and 4) further analysis on the performance on benchmark tasks. The comprehensive results provide additional empirical validation of the performance of the proposed method.
A-E1 The other comparison results for reward model performance
Fig. A.1 shows the comparison between the prediction value by the learned rewards and their ground truths for LEASE, FEWER, and FRESH under Adroit tasks, where the offline algorithm is IQL. The predicted and true rewards are both normalized to $[0,1]$. We randomly sample $500$ data from unlabeled datasets that are not seen in the training stage for evaluation, like Fig. 4. Since the reward has a value in certain positions in Adroit tasks, the figure is in the form of multiple scattered lines in Fig. A.1. As evidenced by the experimental results shown in this figure, the reward model trained under the LEASE framework achieves the smallest variance in predicted reward values, which shows that the data augmentation and selection mechanism can improve the reward model’s performance.
A-E2 The further comparison results between LEASE and URLHF
Considering that it is crucial to include a comparison with URLHF using the same amount of data as LEASE. To further show the superior performance of the proposed method, we compare LEASE to the baseline algorithm URLHF with the same amount of data as LEASE, where the latter method is denoted as $\text{URLHF}^{*}$. Table A.1 shows that the average performance of LEASE is superior to that of the baseline algorithm URLHF using the same amount of data with LEASE. The experimental results demonstrate that the proposed algorithm LEASE achieves a performance improvement of approximately 3.9% compared to $\text{URLHF}^{*}$. This further validates the effectiveness and advantages of the proposed framework and method.
TABLE A.2: Comparison results of offline RL algorithms under the designed framework.
Task Name $\textbf{CQL}^{*}$ $\textbf{IQL}^{*}$ $\textbf{COMBO}^{*}$
walker2d-m $78.4\pm 0.9$ $73.0\pm 2.1$ $71.6\pm 2.4$
walker2d-m-e $98.6\pm 18.1$ $108.1\pm 0.5$ $79.1\pm 1.1$
hopper-m $56.5\pm 0.6$ $56.0\pm 0.5$ $54.8\pm 0.9$
hopper-m-e $6.4\pm 0.8 55$$5.9\pm 1.9$ $54.9\pm 1.1$
halfcheetah-m $43.5\pm 0.1$ $43.0\pm 0.3$ $42.9\pm 0.1$
halfcheetah-m-e $53.2\pm 3.1$ $60.2\pm 1.2$ $73.8\pm 7.0$
Mujoco Average $4.4\pm 4.0 66$$6.4\pm 1.1$ $62.9\pm 2.1$
A-E3 The results for LEASE under model-based RL offline algorithm
To enhance the generalizability of the proposed framework, we extended our experimental validation beyond model-free offline RL approaches (such as CQL and IQL) to include model-based offline RL methodologies. Table A.2 shows the results of COMBO [49] under the designed framework, where * denotes the results of offline RL algorithms under our designed framework. For COMBO hyperparameters, the probability confidence $\kappa_{p}$, and uncertainty variance $\kappa_{\tau}$ are set as $0.85$ and $0.05$ for all tasks, respectively. Note that in our framework, model-based methods do not necessarily perform better than model-free methods. Model-based RL methods focus on how to learn a conservative policy by regularizing $Q$ values or penalizing rewards to alleviate the effects of inaccurate model data. These methods critically depend on the accuracy of the reward model. Consequently, model-based approaches are more sensitive to reward modeling errors than model-free methods, which can hinder the improvement of agent performance. Therefore, model-based offline RL requires a higher accuracy of the reward model than model-free RL.
A-E4 Further analysis for the results on benchmark tasks
Table A.1 shows that LEASE doesn’t achieve better results on all tasks. Compared to URLHF*, LEASE shows significant improvements on several tasks, such as walker2d-m-e, pen-expert, and hammer-expert. For tasks like walker2d-m, hopper-m-e, halfcheetah-m, and door-human, the performance shows slight improvements. It is widely accepted that the current improved RL algorithms struggle to perform well across all tasks. The network architectures and parameters of URLHF and LEASE differ, which could also lead to the performance discrepancy. It can also be observed that in the hopper, halfcheetah, and door environments, the performance improvement of LEASE over URLHF* is minimal, and in some cases, even lower. This is because the accuracy of the reward function in these environments does not have a significant impact on performance. Table A.3 has validated this conclusion. This table shows the performance comparison of the agent in the above three environments using official URLHF code, with different numbers of preference data $100$, $500$, and $1000$.
TABLE A.3: Comparison results between different numbers of preference data.
Number hopper-medium halfcheetah-medium door-expert
$100$ $56.6\pm 2.4$ $43.3\pm 0.2$ $103.3\pm 0.5$
$500$ $55.8\pm 0.1$ $43.4\pm 0.2$ $103.8\pm 1.0$
$1000$ $55.6\pm 2.8$ $43.3\pm 0.2$ $103.6\pm 0.4$